10:00 AM, EAF CONFERENCE ROOM

Chair:   Piet Martens
Minutes: Robert Doney

Present: Joe Covington (MDI), Joe Gurman and Jeff Newmark (EIT),
         Chris StCyr (LASCO), John Raymond and Joe Michels (UVCS),
         Udo Schuehle, Philippe Lemaire, and Tom Moran (SUMER), 
         Andrzej Fludra (CDS), Art Poland and Vicente Domingo 
         (Project Scientists)


1	Review of action items
        - All items closed

2 	Boundary conditions for the coming month
	- MDI Continuous: 1/15 6:50 UT - 1/18 21:50 UT
	- S/C orbit trim, station keeping, and momentum management: 1/14
	- Memory dump 01 Feb 97
	- No TR maintenance
	- Preliminary outlook on March Roll Maneuvers:
	  Roll Maneuver: approximately 36 hours
	  not including: Gyro Calibrations, Momentum Management, 
	  		 and a possible Station Keeping
          MDI has requested special program for March roll, requiring
          22.5 degree steps and dwell after each. SWAN wants to use
          roll for cross-calibration
        - MDI also requests orbit trim and station keeping just prior
          to 3 months continous, i.e. in the first half of April, so
          that continuous program will be interrupted once at most
          for maneuver. Relative timing of trim, roll, and momentum
          management tbd (FOT, SOC's and experiments)

3 	Priorities for the coming month:
	a) Individual instrument plans
	   CDS:    JOP018, Will run a new synoptic study (4 x 2" slit size)
		   which will permit an additional 3 hours of viewing each
		   day by sacrificing some resolution.  Coronal Hole Jets
		   to be run with EIT (no date yet). JOP021 shifted to
                   week 4, and possible run in week 2.

	   SUMER:  JOP022, JOP058 have priority. Other JOPs as listed
                   in calendar. SUMER and UVCS plan collaboration in
                   January, details of which tbd. (More news in Jan.).

	   UVCS:   High Altitude/Coronal Hole Study [Mg(X) at approx 3
		   Solar Radii]

	   LASCO:  Performing Submode 2 transition at the end of this week.
		   This is a c1-c3 cycle with 195 Angstrom EIT exposures,
		   as well as a CME watch (FeX or FeIVX).  LASCO will also
		   try to coordinate a Galileo radio occultation study
                   around 14 January
		   (there is some uncertainty here since Galileo's high
		   gain antenna failed to operate some time ago).  Submode5
		   transition can not be performed before Jan. 20.

	   EIT:	   Will perform a CCD Bakeout during week 1 or 2.  There
		   will also be JOP022 participation, and the last part
		   of January is dedicated to Microjet studies.

	   MDI:    HiRes Study Jan. 15-18 (will try to combine with JOP022);
		   this depends however on what is in the FOV.  Jan. 7-13
		   possible calibration activities

   	b) Calibration plans
           - Update on Intercalibration JOPs:
	       ICAL001 and 002 will be combined into one JOP by Udo Schuehle
               ICAL009 on hold indefinitely, will be rewritten by Joe
           - Schedule for intercalibration activities
               ICAL001 tentatively scheduled for week 4.

   	c) Joint plans

           SOC thanks instrument teams for putting together a good program
           for January in very little time. Also, instrument planner and 
           SOL positions are all filled for the month.

           For detailed January coordinated observing program see:


4	JOPs status  

	- Status of Campaign Catalog, Webform for JOP-leaders' reports:
	   The Campaign database has been redesigned to accept updated 
           information from EOFers.  An email message will be sent out 
           to external JOP leaders with a request to complete the information
           on their past campaigns within the next 2 weeks. Webform will
           contain information on how to make such modifications.
           SOC's will screen inputs to eliminate hoaxes.

5	New JOPs
	- JOP058; No comments

	- SOHO-Yohkoh workshop, March 3-7, 1997
	  Subject: Sept. 25 - Oct. 5 Eruptive Event Observations

	   LASCO: Russ Howard & Chris St. Cyr
    	   SUMER: Jean-Claude Vial (to be confirmed)
	   CDS:   Richard Harrison & Andrzej Fludra
	   EIT:   TBD
	   UVCS:  John Kohl
   	   MDI:   Craig DeForest

Because of the workshop, the 36 hour Roll Maneuver should not take place
during the week of MAR 3-8

* For the ISTP Meeting Jan 7-9 (BLDG 26, rm 205), SoHO instrument
teams should try to get real collaborations with other ISTP missions

* End of January ISTP workshop in Yosemite. ISTP wants to start
producing video there, with results from ISTP missions.

* MDI would like to permanently implement the reduction of VC2 time
  from 10 minutes to 5 minutes at the beginning of NRT during the passes.
  This is the proposed format:

	------------			-------------
	VC2 (5 minutes)			VC2 (5 minutes)
	VC4				VC4
	1 hour VC2			

* Upcoming Meetings: - JAN 17:	Next SWT
                     - JAN 24:	Next SPWG

* Meeting Adjourned: 11:00 am.